OpenStudy (australopithecus):

How do I solve this integral? \[\int\limits_{-\infty}^{0} \frac{x^{2}}{9 + x^{6}}dx\] I have tried trig sub but it is really ugly. I assume by completing the square but the answer is pi/9

OpenStudy (anonymous):

its an improper integral

OpenStudy (australopithecus):

note that I took this integral from \[\int\limits\limits\limits_{-\infty}^{\infty} (x^{2}/(9 + x^{2}))dx\]

OpenStudy (australopithecus):

oh crud I wrote it wrong

OpenStudy (anonymous):

you have to set up the limit as x approaches infinity of that integral from t to 0 instead of infinity to 0

OpenStudy (australopithecus):



OpenStudy (australopithecus):

yeah I know this question is in regards to finding the integral of this function

OpenStudy (turingtest):

you can long-divide, then use partial fractions

OpenStudy (turingtest):

oh that's a six in the bottom?!?

OpenStudy (turingtest):

did I rewrite it correctly?

OpenStudy (blockcolder):

Let u=x^3.

OpenStudy (turingtest):

^yep

OpenStudy (blockcolder):

It will become arctan(something).

OpenStudy (anonymous):

you could make it a product and use int by parts

OpenStudy (turingtest):

that would be a huge mess

OpenStudy (anonymous):

you could derive a quotient rule base integration formula in the same way you get the int by parts one :P

OpenStudy (australopithecus):

the integral is \[\int\limits_{-\infty}^{\infty} x^{2}dx/(9+x^{6})\]

OpenStudy (australopithecus):

trig sub also seems hopeless :\

OpenStudy (blockcolder):

Have you tried letting u=x^3 to get rid of the x^2 on top?

OpenStudy (anonymous):

okay maybe we have to evaluate the riemann sum

OpenStudy (australopithecus):

yeah that seems like the best way to solve it :)